Calculate the molality of 10.94g C6H5OH in 220g of H20
1
Expert's answer
2020-12-10T07:53:15-0500

M = n (C6H5OH) / V раствор

М (C6H5OH) = 94

n (C6H5OH) = m / M = 10,94 / 94 = 0,116 моль

V (C6H5OH) = 0,116 * 22,4 = 2,6л

V (H2O) = 0,220 * 1 = 0,220л

М = 0,116 / 2,82 = 0,04